### v1

# Balance Correlations, Agentic Zeros, and Networks: The Structure of 192 Years of War and Peace

Abstract
Heider’s (1944; 1946; 1958) original P  O  X balance theory predicts human relations based on shared (or differing) perceptions and attitudes between a pair of individuals (P  O) towards an inanimate object X. Social network extensions of his theory have replaced this X with a third individual (Newcomb,, 1953; Cartwright and Harary,, 1956), leading to a plethora of adaptations that have often been inconsistent with Heider and with each other. We present a general model and formal notation for these social network extensions that permit social scientists to be more explicit about their balance theoretic statements. Specifically, we formulate statements as a comparison of two conditional probabilities of a tie, where the conditionals are defined by the 2-path relation Ego  X  Alter:
$P(Ego\stackrel{{\scriptstyle q}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle r}}{{\text{---}}}X\stackrel{{\scriptstyle s}}{{\text{---}}}Alter)\quad>\quad P(Ego\stackrel{{\scriptstyle q}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle r}}{{\text{---}}}X\stackrel{{\scriptstyle s}}{{\text{---}}}Alter))$
where q, r and s, represent different mutually exclusive relations. Given the importance Heider assigned to the role of negative associations, we further identify negative ties as completely separate from non-ties (neutral or zero-valued ties) and consider a signed graph to be a restricted multigraph composed of three mutually exclusive and exhaustive relations: positive ties, negative ties, and zero-ties. We stress that neutrality is the result of a triadic process. Combining these two features into our theoretical frame results in 27 identifiable configurations. Drawing on the work on Transitivity Correlation models, we propose a set of simple descriptive statistics to measure the extent to which evidence for any stipulated balance configuration is present in a network by correlating the presence of a particular tie with the corresponding conditional 2-paths that surround that tie. Finally, we demonstrate how to apply this approach to assess network-level balance in a large data set consisting of friendly vs hostile relations between countries from 1816 to 2007. We find strong evidence particularly for one of the four classic Heiderian balance theory predictions, and virtually no evidence in support of the imbalance predictions. However, we do find stable and surprising evidence that the ’neutral’ ties are important in balancing the relations among nations.
Keywords Balance Theory, Balance Correlations, Network Level Measures
1 Introduction
Balance theory prevails as one of the most widespread theories of social behavior to emerge in the past 70 years. With its roots in psychology Heider, (1958), it is widely used in sociology (e.g., Moody,, 1986; Rawlings and Friedkin,, 2017), political science (e.g., Krasner,, 1985), social psychology (e.g., Newcomb,, 1953), and anthropology (e.g., Hage,, 1979). But, perhaps no field has invested more in balance theory than social networks. Whether these networks are studied by social scientists (e.g., Doreian and Krackhardt,, 2001; Feld,, 1981), statisticians (e.g., Snijders,, 2001; Fienberg et al.,, 1985), or physicists (e.g., Kirkley et al.,, 2019), all appeal to balance theory as an explanation or motivation for actor behavior in their research.
One problem with its ubiquitous presence, however, would be that balance theory has taken on different meanings and forms as it has been adapted by the various branches of the study of human behavior, especially those studying social networks. Different network scholars have reinterpreted balance theory in such different ways that it is almost unrecognizable by those who founded the theory itself.
The purpose of this paper is to provide a formal approach to the study of balance in networks that is precise yet flexible, simple yet sophisticated, and innovative yet hones even more closely to Heider’s original theoretical logic. In a major departure from the dominant research agenda, we look beyond the traditional dichotomous perspective that the only ties of interest are the positive and negative ones. Specifically, we ask the question, what role do agentic zeros (neutral relations) play in the consideration of balanced states in a network? Our approach permits researchers in all fields to better specify and understand balance theory’s implications and predictions under such expanded conditions.
2 Balance in a new theoretical frame
2.1 Background
Heider (1944; 1946; 1958) originally defined balance as a function of the relationships between three entities: a person (P), another person (O), and an inanimate object (X). Each relationship between these three entities was characterized as either positive or negative. If the product of the signs is positive, the triad is deemed to be balanced; if the product is negative, the triad is deemed to be unbalanced. This yields eight combinations of possible tie patterns. The central premise in Heider’s balance theory is that P strives to attain balance due to adapting the relation to O, such that the triad becomes in any of the ’Balanced’-states shown in Table 1.
P  O P  X O  X Balance state
+ + + Balanced
+ - - Balanced
- + - Balanced
- - + Balanced
+ + - Unbalanced
+ - + Unbalanced
- + + Unbalanced
- - - Unbalanced
Table 1: Heider’s P  O  X designation.
Newcomb, (1953) was among the first to apply Heider’s theory exclusively to a triad of people. He specifically suggested that identifying X as a third individual, rather than an object as Heider had done, allowed one to use balance among three people to explain many relational social dynamics. A subsequent important paper in the development of balance theory is Cartwright and Harary (1956), which proposes formal mathematical statement about balance properties of the network as a whole, not just an isolated triple of actors. The triple was considered an unordered set without any special consideration given to one node over another — i.e., it was an unlabeled triple. However, this has obfuscated the focus on the role of “P” as the actor/agent of interest in Heider’s work. Indeed, the labels in “P  O  X” were replaced with far less discriminating labels ${i,j,k}$. This became the standard for one branch of balance theory, the concept and measure of triad census as introduced by Holland and Leinhardt, (1979) and continued to this day (e.g. Dinh et al.,, 2023).
Rapoport, (1963) translated this technical mathematical formulation into simple theoretical statements that inspired subsequent work in this area. He noted that each of these eight configurations had a social interpretation that could be construed to give these balance rules an intuition using concrete, everyday terms like friendship and enemy. He noted that the eight theoretical statements corresponding to Heider’s P  O  X model are (see Table 1):
Balanced Unbalanced
A friend of a friend is a friend. A friend of a friend is an enemy.
A friend of a enemy is an enemy. A friend of an enemy is a friend.
An enemy of an enemy is a friend. An enemy of a friend is a friend.
An enemy of a friend is an enemy. An enemy of an enemy is an enemy.
Table 2: Putting balance theory in words.
While these aforementioned simple Rapaport statements make the Heiderian predictions easy to understand, these also are inherently overly demanding. For example, no one really believes that everytime someone encounters a friend of a friend that she will establish a friendship with that person. Rather, a more reasonable prediction from this first Heiderian principle is that a friend of a friend is more likely to be a friend than if that person is not a friend of a friend. That is, the 2-path relation, friend-of-a-friend, creates a favorable context in which a friendship is more likely to form.
2.2 Balance theory as a set of probability statements
We will throughout this paper replace Heider’s P  O  X terms with corresponding network terms. We start by noting that the only relationship that Heider specifically designated as interpersonal was the relationship between P and O. And the question posed by the model is what relation P would choose to have with O, given the association both have with X. This X provides a contextual element that influences P’s choice. Consistent with the social network literature, we refer to this P  O relationship as an Ego  Alter relationship, wherein Ego is choosing a relationship with Alter. We retain the label “X” to refer to a contextual third party. With this minor amendment, we can formalize the first Heiderian “a friend of a friend is a friend” principle with the following conditional inequality:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)>P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))\hss$ (1)
or, equivalently:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)-P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))>0\hss$ (2)
where $\stackrel{{\scriptstyle f}}{{\text{---}}}$ indicates a friendship tie. Similar conditional inequalities can be made of each of the other three primary Heiderian predictions about balanced relations. In cases of imbalanced relations, such as “a friend of an enemy is a friend,” then the prediction is that the 2-path relation, being a friend of an enemy, should lead to relatively infrequent friendships. Such imbalanced statements would be similarly described as in (1), except that the inequality is reversed:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)<P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))\hss$ (3)
or, equivalently:
$\hskip 36.135ptP(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter)-P(Ego\stackrel{{\scriptstyle f}}{{\text{---}}}Alter\ |\ \neg(Ego\stackrel{{\scriptstyle e}}{{\text{---}}}X\stackrel{{\scriptstyle f}}{{\text{---}}}Alter))<0\ \hss$ (4)
where $\stackrel{{\scriptstyle f}}{{\text{---}}}$ indicates a symmetric friendship tie and $\stackrel{{\scriptstyle e}}{{\text{---}}}$ indicates a symmetric enemy relation.
At the network level, we can frame this prediction as a statement about the relative likelihood that the predicted tie would appear given that it is facing a contextual 2-path relation. That is, in the network as a whole, we can cross-tabulate the number of occurrences of a tie type with the number of occurrences of the specified 2-path relation:
 Is there a 2-path Ego  X  Alter?
1=Yes 0=No
Is there a tie Ego  Alter? 1=Yes A B
0=No C D
Table 3: Cross-tabulation of Ego  Alter ties with 2-path triples, Ego  X  Alter. A, B, C and D are counts of co-occurrences.
If we assume a uniform and random draw from all $\{i,j,k\}$ triples in the network, then we can directly calculate the relevant conditional probabilities from this 2x2 table. The first conditional in Eq. (1) and (2), the probability that a friendship between Ego and Alter exists, given that there is an X who is also a friend of Ego and a friend of Alter, is calculated as $\frac{A}{A+C}$. The second conditional in each equation are calculated as $\frac{B}{B+D}$. Thus, the difference between these two fractions $\frac{A}{A+C}-\frac{B}{B+D}$ represents the extent to which the particular balance theory prediction is true. The larger the difference, the stronger the effect size given the conditional 2-path. If the the relation is unbalanced, such as described in Eq. 3 and 4, then the difference between these two fractions would be negative. The larger the negative value, the stronger the effect of the 2-path in suppressing the existence of the Ego  Alter tie.
Conversely, if $\frac{A}{A+C}-\frac{B}{B+D}$ is positive, then it will also be the case that the difference between the corresponding row percentages will be positive (i.e., $\frac{A}{A+B}-\frac{C}{C+D}>0$). This represents the symmetric claim that if the Ego  Alter tie exists, then the probability of the designated 2-path Ego  X  Alter tie will be greater than if the Ego  Alter tie does not exist. While these differences will not generally equal each other, they will always be of the same sign. Moreover, the size of these differences will correspond to the effect size of the particular Heiderian balance principle.
Putting this altogether, we can describe the direction and size of the balance theory effect by taking the geometric mean of these two differences in probabilities, these two descriptors of balance in the network:
$\sqrt{\left(\frac{A}{A+C}-\frac{B}{B+D}\right)\left(\frac{A}{A+B}-\frac{C}{C+D}\right)}$
Finally, we draw on earlier work Dekker et al., (2019) to note that this geometric mean is equal to the Pearson correlation between the vector of their corresponding Ego  X  Alter 2-path ties for all $\{i,j,k\}$ triples and the corresponding vector of their associated Ego  Alter ties. That is,
$\rho=s\sqrt{\left(\frac{A}{A+C}-\frac{B}{B+D}\right)\left(\frac{A}{A+B}-\frac{C}{C+D}\right)}$ (5)
where:
$s=sgn\left(\frac{A}{A+C}-\frac{B}{B+D}\right).$ (6)
The sign function in 6 differentiates between balanced and unbalanced predictions. If $\frac{A}{A+C}>\frac{B}{B+D}$, then the correlation between the presence of the predicted Ego–Alter tie and its conditional 2-path will be positive; thus, the positive root of this square root function is used, indicating that the Ego  Alter tie probability is enhanced in the presence of the Ego  X  Alter 2-path. If $\frac{A}{A+C}<\frac{B}{B+D}$, then this correlation will be negative; thus, the negative root of this square root function is used, indicating that the Ego  Alter tie probability is suppressed in the presence of the Ego  X  Alter 2-path.
In other words, the Pearson product-moment correlation between observed Ego  Alter ties and their associated Ego  X  Alter 2-path ties directly measures the extent to which balance (or unbalance) is observed in the network.
2.3 The Critical Role of (Agentic) Non-ties
Another restriction in Heider’s model is that it assumes all relations are either positive or negative, not neutral. On the other hand, as Cartwright and Harary noted (1956) very early on, in reality many people are going to have neutral sentiments (neither positive nor negative) about each other. While such non-ties are often seen as part of any natural network, they are often treated as “missing data” and are excluded from any predictions (Cartwright & Harary, 1956). Or, by some conventions, as noted by Rawlings and Friedkin (2017, p. 515), “… the absence of a positive sentiment implies the presence of a negative sentiment,” and as such a lack of a tie may be deemed to represent a negative tie. We disagree completely.
In a major departure from this tradition, we explicitly consider neutral relations as agentic or social choices, just as much as positive and negative relations. There are many reasons that neutral or non-relationship ties should be considered within a balance theory context. For example, in a highly polarized and conflict-ridden environment, Ego may have positive sentiments toward X, but Ego also observes that Alter and X have an intense dislike for each other. One way of escaping this uncomfortable position is to establish a neutral or non-relationship with Alter, rather than taking sides in the dispute. That is, Ego may take the “Switzerland” option of staying neutral in a spat, in order to preserve his options to work with Alter in the future.
Strategic neutrality is not the only reason for considering neutral relations in balance theory. Within the blockmodeling perspective it has been noted frequently how zero blocks (groups of people who are discouraged from relating to each other or to other groups) can be found as a way to preserve social order (White, Boorman & Breiger, 1976; Pachucki & Breiger, 2010). The cost of forming ties under such conditions could be substantial in a social situation. Understanding how these non-ties play a role in balancing triadic relations is explicitly part of our model.
Thus, Ego may develop a neutral attitude towards Alter. Moreover, Ego’s observation that X may have a neutral relation with either Ego or Alter (or both) could affect Ego’s calculations in how to consider her tie with Alter. We therefore maintain that any prediction stemming from balance theory about a real network must recognize this inherent character of the network and include positive, negative and neutral ties as part of the balance processes.
2.4 Triad Signature
Our claim, then, is that all signed graphs can be decomposed into three (sentiment) subnetworks: Positive ties, Negative ties, and Zero ties (neutral or non-ties). These three types of relations define a restricted multigraph on any network $\mathcal{G}$:
  • {P} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ has a positive tie with $j$;
  • {N} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ has a negative tie with $j$;
  • {Z} :=assign:=:= the set of ($i,j$) unordered pairs designating all ties where $i$ does not have a positive tie nor a negative tie with $j$, which we indicate as a neutral or “zero” tie.
We ignore self-ties ($i,i$) and hence these will not be in these three sets.
By convention, we represent these three sets in square matrix form as follows:
  • $\textbf{P}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{P}\}\\ 0,&\text{otherwise}\end{cases}$
  • $\textbf{N}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{N}\}\\ 0,&\text{otherwise}\end{cases}$
  • $\textbf{Z}_{ij}=\begin{cases}1,&\text{if ($i,j$)}\ \in\{\textbf{Z}\}\\ 0,&\text{otherwise.}\end{cases}$
These sets are, by definition, mutually exclusive and exhaustive in the network; thus, their sum as well as their union forms a complete and simple graph on the full set of nodes in $\mathcal{G}$. These three networks (P, N and Z) henceforth will be referred to as the set of primitive relations in $\mathcal{G}$.
Given the three possible relations in any ($i,j$) pair, then there are $3\times 3=9$ possible configurations in any 2-path $i\text{-}k\text{-}j$. For each 2-path configuration, there are three relations between $i$ and $j$ that complete the triple, yielding $273^{3}=$ 27 configurations possible. Each of these may be designated with a 3-tuple identifier. We will use the following rules to identify each unique possible configuration:
Given any ordered triple of actors in graph $\mathcal{G}$, we designate three roles, one ascribed to each of the three actors: Ego, the focal person to whom the balance consideration is paramount; Alter, the person whom Ego is potentially establishing a relationship with; and Context, the person with whom both Ego and Alter have a relation that defines the context that determines whether the $i\text{-}j$ relation is in balance or not. We will heretofore abbreviate these role references as “E” for Ego, “A” for Alter, and we retain Heider’s “X” for the Contextual third party. Any labeled triple, then, has three tie values, designated by the three pairs of the three nodes. Each pair will be accorded a value corresponding to one of the three types of ties (Positive, Negative, or Zero — from hereon out referred to as $p$ ties, $n$ ties and $z$ ties, respectively). Using this notation, any and all configurations of ties amongst any three nodes may be designated by assigning the letter $p$, $n$ or $z$ to each of the three pairs. The first letter will designate the relation between Ego and Alter; the second will designate the relation between Ego and X; and the third letter will designate the relation between Alter and X (see Table 4).
Position Tie Designation
1 $E\stackrel{{\scriptstyle q}}{{\text{---}}}A$ (Ego with Alter)
2 $E\stackrel{{\scriptstyle r}}{{\text{---}}}X$ (Ego with X)
3 $A\stackrel{{\scriptstyle s}}{{\text{---}}}X$ (Alter with X)
Table 4: Triad Signature Position Designation: The triad signature is a 3-tuple. Each edge, $q,r,$ and $s$, designates an undirected tie of type $p,n,$ or $z$ among the triple labeled Ego, Alter and X.
For reasons that will become clear later, we will separate the first letter in this 3-tuple from the remaining two letters with a period. Thus, for example, the 3-tuple <p.nz> would indicate that Ego has a positive tie with Alter, Ego has a negative tie with X, and Alter has a zero tie (non-tie) with X. Heider’s “the enemy of an enemy is a friend” prediction would be represented as <p.nn>. These three letters, in the designated order, provide a triad signature, which permits the unique designation of each of the $273^{3}\ =$ 27 possible configurations that Ego might face.
While any ordering of these letters is possible, we have chosen this arbitrary ordering because it allows one to see some of the principles behind Heider’s theoretical predictions. In fact, the consistent application of this order of these three elemental relations in the signature provide three important features: 1) The order identifies a specific structural configuration, one of the possible 27; 2) The order represents specific and formal predictive statements from balance theory; and 3) The order naturally leads to a precise calculation of the balance correlation measure of the network-level of balance specifically designated by the signature. We will discuss each of these features in turn below.
2.4.1 Identifying the 27 Triad Signatures
The 27 possible labeled configurations are provided in Figure 1. Positive ties (p) are represented in black; negative ties (n) are represented in dashed red lines; zero-ties (z) or neutral ties are blank, and are indicated with a letter “z” to emphasize they are treated as ties. Node labels refer to the Ego (E), Alter (A), and Context (X) roles in each triad. As we will discuss later on, at the network level some of these triads will be obfuscated. Nonetheless, in order to keep the theoretical stories clear, it is still important to retain all 27 triad signatures.
As an interesting extension of this last point, we note that many of these labeled configurations are isomorphic, yet their theoretical claims are still distinct. For example, <p.pz> is isomorphic with <z.pp>, but each is a different claim on the experience that Ego faces. In the former case, Ego is the central character, the bridge between two friends who are neutral toward each other. In the <z.pp> case, however, Ego is neutral to Alter, even though his friend (X) likes Alter. How each of these two configurations may play out in Ego’s experience from a balance theory perspective could be quite different, depending on the options Ego experiences. In the <z.pp> case, Ego could actively close the triple with a positive tie to Alter, creating a <p.pp> triple, which would be attractively balanced in Heider’s view. But, in the <p.pz> case, Ego has less agency, since the neutral tie is between Alter and X. Thus, these two isomorphic structures represent entirely different balancing experiences and options from Ego’s viewpoint.
Figure 1: The 27 Triad Signatures.
Note: Each signature is given by the three letters, any combination of p, n and z, in angle brackets. A black solid line indicates a positive tie; a red dashed line indicates a negative tie; no line indicates a zero-tie (a neutral or non-tie). Node roles of Ego, Alter, and Contextual third party are represented by the letters E, A and X, respectively.
1
2.4.2 Triad Signature as a Predictive Statement
Each triadic signature also operates as a unique theoretical predictive statement regarding balance processes. To see how this is so, we return to the predictive statement that Heiderian balance can make: If balance theory holds, then Ego should be motivated to create balance in its ties and to alter relations that create unbalanced situations. In a triadic context, the probability that Ego will actively choose to engage in a tie (either p, n or z) with Alter will depend on the 2-path relation through X. In the present model, the first letter in the triad signature represents this choice faced by Ego. This separate role is visually enhanced by the placement of a period between it and the latter two letters in the signature.
To illustrate this feature, we begin with a list of all eight classic Heiderian balance-theory statements, along with our triadic signature notation in Table (5). The core prediction in Heider’s model is that if a triple is unbalanced, then this creates psychological tension and discomfort. Thus, an unbalanced triple should not occur, or if it does occur, an attempt will be made by Ego to balance it by reconsidering her relations with Alter to make it balanced. On the other hand, if a triple is balanced, it will induce a sense of equanimity on Ego’s part and is more likely to be stable and remain in place.
Balanced signature Unbalanced signature
A friend of a friend is a friend. <p.pp> A friend of a friend is an enemy. <n.pp>
A friend of a enemy is an enemy. <n.np> A friend of an enemy is a friend. <p.np>
An enemy of an enemy is a friend. <p.nn> An enemy of a friend is a friend. <p.pn>
An enemy of a friend is an enemy. <n.pn> An enemy of an enemy is an enemy. <n.nn>
Table 5: Rapaport descriptions and their corresponding Triad Signature representations.
Viewed in this way, each balanced signature easily translates to a corresponding prediction from balance theory. Consider the following decomposition of the triadic elements.
Given any set of relational ties ($q$, $r$, and $s$) among Ego, Alter and X, such that Ego has a $q$-tie with Alter, Ego has an $r$-tie with X, and Alter has an $s$-tie with X, the signature of this triad <q.rs> can be decomposed into two parts: the $q$-tie, and the 2-path $rs$-tie. Now all balance theory statements can be framed by treating the $q$-tie as a dependent variable, predicted by the contextual 2-path, the $rs$ component (i.e., $P(q)=f(rs)$).
As a general principle, if the triple <q.rs> is balanced, then the prediction is:
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad>\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss$ (7)
To illustrate, the Heiderian balance prediction that “the enemy of an enemy is a friend” can be restated more formally as “the probability that a specific Alter will be a friend is enhanced if I have an enemy who is also the enemy of this Alter.” This statement immediately corresponds to and is represented by the configuration <p.nn>, where p is the dependent variable and $nn$ is the 2-path conditional. The prediction is that $p$ is more likely to occur given the presence of the 2-path $nn$ than if $nn$ is not present.
$P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X)\quad<\quad P(E\stackrel{{\scriptstyle q}}{{\text{---}}}A\ |\ \neg(E\stackrel{{\scriptstyle r}}{{\text{---}}}X\wedge A\stackrel{{\scriptstyle s}}{{\text{---}}}X))\hss$ (8)
Take as an example the second unbalanced statement in Table 2, “A friend of an enemy is a friend,” represented by the configuration <p.np>. Since it is theoretically unbalanced, it is predicted to create an uncomfortable situation for Ego. Given that the 2-path relation “a friend of an enemy” exists between Ego and Alter, then it is predicted that Ego will not, or at least be less likely to, establish a positive friendship tie to Alter. The specific prediction, then, is that a p-tie is less likely to occur given the presence of the 2-path np than if np is not present.
All balance theory statements can be framed this way. Evidence for or against these predictions, stemming directly from triad signatures, can be brought to bear on each statement to see which specific predictions hold up under scrutiny. While this sounds straightforward, we will soon see that the issue of balance is considerably more complicated than this simple interpretation.
2.4.3 Triad Signature-induced Measure of Network Balance: The Balance Correlation
The Embeddedness Principle. Before we introduce this third, and most critical, function of the Triad Signature, we note one final difference between Heider’s original theoretical formulation and applications to social network analysis. Heider himself focused on isolated and ideational triples of entities and made predictions based solely on these three entities. In a network context, however, it is important to recognize that no labeled Ego  X  Alter combination occurs in isolation but rather is embedded in a complex structure of triples. In fact, changes in one tie can imply balance changes elsewhere in the network. Consider the case depicted in Figure 2. The three nodes in the middle of the graph (Ego, 1, and 2), constitute an unbalanced triple: Ego has a positive tie with Node 1, and Ego also has a positive tie with Node 2; but, Nodes 1 and 2 have a negative tie. Consider the specific ordered triple where node 1 is Alter and node 2 is the contextual X. Ego is faced with a dilemma. In considering whether to change her relationship with Alter (node 1), Ego finds herself embedded in several critical triples with node 1 as an Alter ((Ego,1,3), (Ego,1,4), (Ego,1,5)). Each of these triples has the signature of <p.pp>, and therefore these are balanced. If Ego were to balance the (Ego,1,2) triple by switching her relationship with node 1 from <p> to <n>, then that would distort the balance that currently exists in these three other <p.pp> triples. In Heiderian terms, the balance cost would far outweigh the balance benefits for switching.
Figure 2: Example of a Network Illustrating the Embeddedness Principle
Note: The black edges represent positive ties; the red dashed edges represent negative ties. The focal labeled triple of interest here is {Ego,1,2}, which is unbalanced (<p.pn>). Since the Ego-2 tie is embedded in more unbalanced triples than the Ego-1 tie is, Ego is more likely to change her relationship with node 2 as an Alter than node 1 as an Alter in trying to resolve this unbalance.
On the other hand, consider the alternate ordered triple where Ego is now taking stock of her relationship with 2 as the Alter, in light of node 1’s new role as Context (X). With node 2 as Alter, Ego also has to consider that there are, again, several contexts ((Ego,2,6), (Ego,2,7), and (Ego,2,8)), to take into account, not just the triple with node 1 (Ego,2,1). In this case, however, all the pertinent triples are currently unbalanced with a <p.pn> structure. If Ego were to switch her relation to Alter (node 2) from <p> to <n>, then all four of these ordered triples would suddenly become a balanced triple <n.pn>. Therefore, in this case, it is easy to see that the balance forces greatly favor Ego’s choice of changing her tie to Node 2 from positive to negative to fix the unbalance in her embedded network.
Not all complex contexts are as easily resolved as in this artificial example. What needs to be underscored here, though, is that in a network context, the Ego  Alter pair does not have one sole X to consider in any balance calculation. And, the larger the number of contextual 2-paths that constitute pre-conditions for balance, the greater the nudging force on the Ego  Alter pair to make a relationship that obeys the rules of balance. Indeed, each Ego  Alter pair faces (n-2) triples of some configuration; some of these may be balanced, some unbalanced, and perhaps some undecided. The Embeddedness Principle simply states that all of these triples must be considered in assessing what the balance forces are on any Ego  Alter pair.
Taking this Embeddedness Principle into account, we now make Heider’s prediction more concrete. Since social networks provide many observations, even around a specific dyad, this affords us the opportunity to turn the prediction into a measurable probabilistic statement, rather than simply “the network is balanced” or “the network is unbalanced”, as Cartwright and Harary, (1956) originally suggested. We propose that balance statements are better framed with words like “tends to” or “is more likely to”. Consider the “friend of a friend is a friend” prediction. Rather than stating “In every case, a friend of a friend is a friend”, we argue that a better, empirically testable extension of this claim would be “A friend of a friend will tend to be a friend.” And the more friend-of-friend contexts one is embedded in, the more likely Ego will choose Alter as a friend, assuming balance theory is correct. This insight will help us explore Heider’s model more precisely.
One of the basic predictions stemming from balance theory is that, if individuals are prone to attaining and maintaining balance, then at any given time we should find that balanced relations are more prevalent in a network. That is, if balance theory actually works to influence overall structures, then we should find that the 2-paths that exist in a network should be associated with a higher probability of the balance-confirming direct tie from Ego to Alter. As we observed in equation 1, the more a balance-based 2-path conditional exists around any particular Ego  Alter pair, the higher the probability we would observe the predicted direct tie from Ego to Alter.
A key advantage of formalizing Heiderian predictions using triad signatures is that they make it easy to extend this correlational measure to include this embeddedness principle. Dekker, et al. (2019) created a correlation coefficient, the Transitivity Correlation, that assesses the extent to which the 2-path pre-transitive condition in a digraph (${i}\text{-}{k}\text{-}{j}$) is correlated with the presence of a direct tie ${i}\text{-}{j}$. Their solution to this problem was simply to correlate the off-diagonal elements in the matrix form of $\mathcal{G}$ with the corresponding elements in matrix $\mathcal{G}^{2}$. Since $\mathcal{G}^{2}$ contains the number of directed 2-path ties of the form $i\text{-}k\text{-}j$ for every $\langle i,j\rangle$ dyad, this point biserial correlation indicates the extent to which there is an overall tendency in the digraph $\mathcal{G}$ to close a directed 2-path with a directed tie. If the correlation is strong and positive, that would indicate that the tendency towards transitive closure is increased as a function of the number of directed 2-path paths surrounding any dyad. A negative correlation would indicate that the more directed 2-path paths that surround a dyad, the less likely it is that $i$ will be directly tied to $j$, indicating that $\mathcal{G}$ has an intransitivity propensity (such as one would find in a star-shaped graph). A zero correlation would indicate that, controlling for the density in $\mathcal{G}$ and the number of directed 2-paths in $\mathcal{G}$, there is no overall propensity towards transitivity in $\mathcal{G}$.
We can use this same approach to look at balance in a graph. While the solution is not as simple as correlating $\mathcal{G}$ with $\mathcal{G}^{2}$, we can still create a matrix of the number of contextual 2-paths that would be specified by the balance theory prediction; and we can just as easily create another matrix of ties that would be predicted by this same balance theory statement. If these matrices are positively correlated, then that would constitute evidence that the specific balance predictions tend to hold in $\mathcal{G}$.
This approach requires calculating for any given signed graph, $\mathcal{M}$, two matrices, $\textbf{M}_{1}$ and $\textbf{M}_{2}$, where $\textbf{M}_{1}$ contains the predicted Ego  Alter relation, and $\textbf{M}_{2}$ contains the number of conditional 2-paths around each dyad that are being tested. These two matrices are straightforward derivatives of the signature notation we have used thus far. In the general case, where the triad signature describing the predicted balance relation is given by the 3-tuple <q.rs>, then
${\textbf{M}_{1}}=\textbf{ Q},\ {\textbf{M}_{2}}=\textbf{RS}$
where $\bf RS$ is the inner product of the relation $\bf R$ and the relation $\bf S$.
A balance correlation, then, is simply the product-moment correlation between off-diagonal elements in $\textbf{M}_{1}$ with the corresponding elements in $\textbf{M}_{2}$. Moreover, as just demonstrated, these correlations are easily generated from the signatures themselves. For example, the <n.pn> signature generates a balance statement that negative n-ties are induced from an abundance of 2-path pn-ties; and the balance correlation is $\rho(\bf N,PN)$, where N is the matrix of n-ties, and PN is the compound word (White, Boorman & Breiger, 1976) created by the matrix product of the P matrix and the N matrix of ties .
Balance theory predicts a positive correlation on balanced signatures and a negative correlation on unbalanced signatures. Whether a network is balanced overall, in this view, depends on which of the possible balanced or unbalanced signatures you are looking at. For example, within the same network, there could be a strong positive balance correlation for the predictive <p.nn> statement but a negative correlation for the predictive <n.pn>. Or, perhaps we observe a positive correlation for <p.pp> but also a positive correlation for <p.pn> (which, according to balance theory, should be negative). That is, balance theory may hold for some configurations but not others.
Phenotypical Twins. As we have already noted above, there are $273^{3}=$ 27 triad signatures. And, each of these signatures doubles as a possible prediction and generates a correlation that permits us to evaluate this prediction. It should be recognized, while these predictive statements and measured correlations can be generated for any of the 27 signatures, many of these signatures do not correspond to anything that Heider would claim actually stems from balance theory. Yet, it is still possible that these empirical correlations might yield insights into structural patterns that could lead to new theoretical extensions, akin to what balance theory has produced over the past half century (e.g., Holland and Leinhardt,, 1979; Davis,, 1979). A thorough exploration of such patterns in any dataset would result in a 9 $\times$ 3 correlation matrix, where each of the three dependent relations (Q, R, and S) is potentially predicted by each of the 9 possible 2-path relations.
While such an exploration sounds inviting, an important caveat is required here. Each of these signatures represents a different theoretical claim/prediction, but it is not the case that each is empirically distinguishable from all other predictions at the aggregated network level. For example, consider the balanced triad signature <n.np>, Ego sends a negative tie to Alter and a negative tie to X; Alter sends a positive tie to X. The prediction represented by this signature is that, conditioned on Ego disliking a third party (X) and Alter likes that same third party, this promotes Ego to send a negative tie to Alter. In assessing this statement, the correlation takes into account how many of these particular conditional 2-paths there are between Ego and Alter.
Now consider a second configuration/prediction, also balanced, represented by the triad signature: <n.pn>. Here the primary difference is that the conditional 2-path represents Ego as having a positive relationship with X and Alter has a negative relationship with X (rather than vice versa). While the balance prediction is the same, the theoretical process is different. Indeed, in the first <n.pn> case, at least Ego has one positive tie, which she may wish to keep. In the second case, Ego is surrounded by negative ties, and perhaps is less committed to maintaining them.
These two predictions, even though both represent balanced states, are quite different from one another on the surface. Yet, while the theoretical mechanism for balance and adjustment might be arguably dissimilar in these two cases, a closer look reveals that they are empirically indistinguishable at the network level.
Consider the particular triple in the first case, <n.np>, where Ego sends a negative tie to Alter. We found in that particular Ego  Alter pair that there existed a certain number of 2-paths of the form <nz>. Let’s call that number $k_{nz}$. It is also the case that for every Ego  Alter pair where the relationship is <n> (negative), there is exactly one corresponding pair where Ego and Alter switch roles as Ego becomes Alter and Alter becomes Ego. For each original Ego  Alter pair, the mirrored Alter-Ego pair will also have the same set of $k_{nz}$ 2-paths, except when they switch roles the <nz> 2-paths become <zn> 2-paths. Thus, $k_{nz}$ for the original Ego  Alter pair will equal $k_{zn}$ for the mirrored Alter-Ego pair. This equality will repeat for each original Ego  Alter pair. These equivalent and commensurate frequency counts of dyads and 2-paths will be repeated for every Ego  Alter pair in the multigraph. The result is that the correlation between N and NP will equal the correlation between N and PN. They will necessarily be identical, even though they theoretically describe two different mechanisms.
We claim that these two configurations, these two triad signatures <n.pn> and <n.np> are behaviorally indistinguishable, or phenotypical twins , in that no evidence can be observed in any network that would allow one to know which of the two theoretical mechanisms is a better explanation of the tendencies in the network toward tie formation or maintenance. While their genotypes are different (they have different fundamental signatures and theoretical underpinnings), their observable manifestations and influences, at least as indicated by exact correlations between the predicted relation and the conditional 2-paths, are not.
Not all signatures have phenotypical twins (hereafter referred to simply as “twins”). Indeed, of the 27 signature triads, nine have no twin. But the remaining 18 do have exactly one twin. It is important to know which signatures have a twin, and what the twin’s signature is, because it will be impossible to tell which theoretical explanation deserves credit for the observed strength in predictive power as indicated by the balance correlation. It will call for a mandatory dose of humility on all our parts.
We provide the following theorem for uncovering twins in triad signatures (a general proof of this theorem on multi-digraphs is provided in the Appendix):
Theorem 2.1.
For any triad signature <q.rs>, if it has a twin, it will consist of the signature <q.sr>. If and only if s=r (that is, s and r are the same relation), then <q.rs> will not have a twin.
Thus, <n.pn> and <n.np> are twins. Note, however, that <n.pn> and <p.nn> are not twins, even though they are isomorphic in their graph structures  [^1] It is important to note that non-twin signatures can, in any given graph, have identical balance correlations. There may be several balance correlations equal to 0, for example, indicating that there is no tendency toward balance or unbalance as predicted in those signatures. The critical issue with twin signatures is that their Balance correlations will be equal by construction in any and all possible graphs of any size composed of any arrangement of ¡p¿, ¡n¿ and ¡z¿ ties.. As we will see below in our analysis, isomorphic triads but which are not twin signatures, can have dramatically different balance correlations. Indeed, in some cases, we will observe isomorphic triadic signatures that will reveal opposite signs; that is, one signature will exhibit balance while its isomorphic counterpart will be unbalanced.
3 Empirical Example
3.1 Data
To demonstrate the usefulness and flexibility of this model, we will use the Correlates of War (CoW) data. These data are the result of an ambitious project started in 1969 by a political scientist at the University of Michigan, David Singer (see Singer,, 1988; Bremer et al.,, 2022, for a current and exhaustive review). This data set includes a comprehensive accounting of both formal defense alliances and military acts of aggression that occurred each year between all pairs of countries from 1816 to 2007. We considered any formal defense alliance between the pair of countries as a positive tie. Acts of aggression recorded in the data set included many categories, ranging from threats of military action against another country to all out wars or invasions (see Table 6 taken from Palmeri et al., 2020). We coded any aggressive act by one country against another as a negative tie. In the rare case where both a defense alliance and an act of aggression occurred within the same year, we coded that tie as a negative tie for that year.
As detailed in Palmeri et al., (2020), there were 22 CoW codes regarding negative ties between nations. They clustered these categories into five buckets, each representing a level of hostility between nations, where 1 represented no military action and 5 represented a full scale war. Instances of no military action (level 1 per their coding scheme) were not coded as a negative tie in our analysis. The remaining 21 codes, however, were deemed a negative tie in this paper. See Table 6 for a full listing of these codes.
Table 6: CoW Codes for Negative Ties between Nations.
Threat to use force [2] Show of force [3] Occupation of territory [4] Begin interstate war [5]
Threat to blockade [2] Alert [3] Seizure [4] Join interstate war [5]
Threat to occupy territory [2] Nuclear alert [3] Attack [4]
Threat to declare war [2] Mobilization [3] Clash [4]
Threat to use CBR weapons [2] Fortify border [3] Declaration of war [4]
Threat to join war [2] Border violation [3] Use of CBR weapons [4]
Blockade [4]
Note: The above 21 coding categories used in the CoW data are organized into severity buckets, labeled [2] through [5], where [5] is the most severe type of event (see Palmeri et al., (2020) for details). We left out the category ‘No militarized action [1]’ as this did not reasonably indicate a negative tie.
3.2 Testing Balance Correlations
Testing a balance correlation raises the question of which null distribution, against which we would compare the observed statistic, is pertinent.
We rely on a Monte Carlo-simulation to generate a sample of random networks from a data generating process (DGP) assumed for the population. This DGP preserves the marginal distribution of negative, positive and zero ties observed in the data for each year. For example, if a country has a defense alliance with 12 other countries, and engages in an act of aggression with 5 other countries, then the DGP would generate sample networks that would preserve this expected number of positive and negative ties with other countries. In addition, our null-generating process preserves the observed covariance in distribution of negative, positive and zero ties in each year. This constitutes a conservative test, one that restricts the sampling space support taking the empirical marginals of the data matrices as expected degree. This degree based approach is more conservative than for example an Erdös-Renyi random network approach, because the latter would allow a wider range of possible random networks to be considered. In our data set of country ties, we assume that there is diversity between countries in their capacity to engage in positive and negative ties. Note that this doesn’t imply countries with less capacity are not making active choices. Their choice sets differ, but we assume they can and do consider all possible ties explicitly or implicitly. This assumption supports our choice to base the test on degree expectations, as it includes the investment distribution all countries make.
A further issue to consider is the fact that the randomly generated matrices are mutually exclusive in their elements, i.e. the sum over the primitive matrices must be exactly one. The procedure we take is that for each element in the network we sample one of three values $\{-1,0,1\}$, with probability based on expected row and column degrees of each of the N, Z, and P networks, respectively. This process converges in the limit to observed degree distribution of the empirical data set.
This DGP approach is based on the observed data for each given year. Thus, the set of possible null graphs can be seriously constrained, and these constraints can change for each balance correlation for each year. The resulting distribution of null graphs may exhibit unfamiliar shapes. For example, in Figure 3 we see the distribution of balance correlations for the set of <pp> 2-path balance correlations (<p.pp>, <n.pp>, and <z.pp>). The <p.pp> balance correlations are mostly positive, ranging from just below zero to just over .5. The <z.pp> balance correlations, however, are mostly negative, ranging from -.4 to about .1. The distribution for the <n.pp> balance correlations, however, is much more peaked, with a pronounced mode of about -.03 but a range from just below that up to +.18. The shape of each of these distributions depends on the number of countries in the data set that year, the density of each type of tie, the degree distributions of each type of tie, and the covariances among these degree distributions. These peculiarities are the reason that an appropriate DGP was used to test each balance correlation separately for each year.
Figure 3: Example of Distribution of Balance Correlations under the Null
Balance Correlation Note: In this example, one thousand null graphs were generated from a model that simulated networks that had expected p, n, and z degrees equal to those observed in 1916. The observed densities of the positive, negative and zero ties were .023, .036, and .940, respectively. This histogram displays the distribution of three balance correlations (<p.pp>, <n.pp>, and <z.pp>) calculated from these null graphs. The blue-shaded part of the histogam, mostly on the right side of the chart, displays the frequency distribution of the <p.pp> correlations. The yellow-shaded part, mostly on the left side of the chart, displays the frequency distribution of the <z.pp> correlations. The red-shaded portion of the histogram, the more peaked distribution in the middle of the chart, displays the frequency distribution of the <n.pp> correlations. The solid vertical colored lines represent the corresponding balance correlations for the observed data for that year. The observed balance correlations for both <p.pp> (the vertical blue line) and <z.pp> (the vertical yellow line) were outside the null region, suggesting they were statistically significant coefficients. The observed balance correlation <n.pp> (the vertical red line), however, was within the null distribution region and therefore not statistically significant.
3.3 Results
Before we present the primary balance results, it is important to comment on the context for these data. Overt acts of aggression (our source of negative ties) vary over the years for which we had data (1816 to 2007). In a few of the earlier years, there were no acts of aggression at all. For such years, we do not, indeed cannot, compute a balance correlation. Also, the number of affirmative defense agreements between nations varies considerably over this period of history. And, of course, this means the proportion of zero ties will also vary over this time.
Figure 4: Density of $p$, $n$ and $z$ Ties per Year
Note: The three dotted plots show how the density of the positive, negative and zero ties varied over the 192 years studied. The blue dots indicate the density of positive ties, the red dots indicate the density of negative ties, and the black dots indicate the density of the zero ties. These densities within any given year by definition sum to 1.0. Of special note is the period of time between 1867 and 1936, the LOA years, during which there was a precipitous decline in the density of positive ties (alliances) between nations.
In Figure 4, we plot the density of ties of each of the three types over this time period. First and foremost, it is clear that Z-ties are the most prevalent, with densities ranging from .75 to .99 over these 192 years. Negative ties are usually (but not always) less frequent than positive ties, with spikes in density most notably during WWI and WWII. Perhaps most interesting is the pattern of positive ties. In general, the density of international relations characterized by formal defense alliances of some type varied between almost none (1909) to almost 20% (at the conclusion of WWII). What is striking in this plot is that these alliances were most frequent before 1867 and after 1936. In between these two years, the density of inter-nation defense agreements dropped considerably. 1867 was the beginning of the dissolution of the Austro-Hungarian empire; 1936 was the onset of WWII with the re-militarization of the Rhineland. We will call this period the “Lack of Alliances” (or LOA) period. We use it here to refer to this specific span of years (1867 to 1936), one that we found to be critical in our exploration of balance, given the regularity with which patterns of balance were interrupted during this period. It can be characterized as a period not of relative calm as much as a period of disorder, where alliances were either not sought or not lasting.
We organize the presentation of the balance correlations into families. A family is defined by the 2-path conditional. For example, the <pn> family consists of the triad signatures <p.pn>, <n.pn>, and <z.pn>.
Thus, we present three results for each of six families (<pp>, <pn>, <pz>, <nn>, <nz>, and <zz>) below. We leave out the phenotypical twins (<np>, <zp> and <zn>) from our presentation of the results below since they would be redundant with <pn>, <pz> and <nz>, respectively.
In each of the following figures, we plot the balance correlations for each year across the 192 years for which we have data (1816 to 2007). In addition, we show the 95% interval from the DGP simulated null distribution in gray. Thus, points representing balance correlations outside that gray region can be considered statistically significant in that they would not be expected to appear under the null model. These significant balance correlations are indicated with a red boundary around the blue dot. Balance correlations that are within the bounds of the null model, thus not deemed statistically significant, are indicated with a small blue dot only.
Figure 5: PP Family of Balance Correlations
3.3.1 PP Family
Figure (5) contains the three plots of the <pp> family of balance correlations for <p.pp>, <n.pp>, and <z.pp>. In the first plot, we examine the pattern of the classic balance claim that a “friend of a friend is a friend.” We see that the <p.pp> balance correlations For the most part are persistently strong, ranging mostly from .8 to .95, providing strong evidence for this balance condition. The primary exception to this lies within the LOA period we identified earlier. During this time, we can see that the country level behavior in the world tends to result in less balance on a global level. It is clear, also, that the low number of positive alliances gives rise to a wider band of balance correlations under the null model. This suggests system wide turmoil.
Remarkably, there is much less evidence for a persistent pattern of unbalance represented by the <n.pp> triad signature. According the classic balance theory, a friend of a friend should not be an enemy, or at least is less likely to be an enemy. If this principle were true, then this should yield a strongly negative correlation. Indeed, there is little evidence that this condition lowers the probability of a negative tie between the two nations; these correlations hover around zero for almost the entire observed period.
While negative ties are not suppressed in the face of 2-path alliance relations, zero ties are clearly so. The null distribution (the gray region in the plot) suggests that the structure of the distribution of positive, negative and zero ties makes a randomly generated balance correlation more likely to be negative, it is still the case that zero ties are much less likely to appear than can be attributed to chance under this null model. Again, the primary exception to this general finding occurs during the LOA period. It appears for the <pp> family, countries are likely to form alliances with other countries where they have many common mutual alliances with third party countries. And, while they may or may not engage in a hostile stance against these countries, they appear to avoid a neutral or non-tie relation with them. That is, they are forced to either fish or cut bait, to choose sides.
Figure 6: PN (and Twin NP) Family of Balance Correlations
3.3.2 PN (and Twin NP) Family
Figure 6, showing the three plots of balance correlations in the PN family, provides an interesting extension of the results in the prior figure. While the balance correlations for <p.pn>, the first panel in Figure 6, show a predominent tendency towards correlations near zero up to the years just prior to WWII, the period subsequent to WWII contain mostly significantly positive correlations. This is true despite the fact that <p.pn> configuration theoretically represents an unbalanced condition: One should not consider an enemy of a friend (or its twin condition, a friend of an enemy) a friend. Yet it seems to be so, and steadily so since the late 1950’s.
On the other hand, the second plot in Figure 6 shows reasonable evidence for balance in that for a majority of years the <n.pn> configuration is also positive and significant. And the strength of these findings, as represented by the size of the correlations, exceeds the size of the balance-disconfirming correlations in the <p.pn> plot. That is, during much of this period of time we examined, the propensity to engage in a hostile act against a country who is a friend of an enemy is substantially greater than if they are not a friend of an enemy. The fact that both of these conditions (<p.pn> and <n.pn>) exist at the same time speaks to the complexity of international relations. To presume that only one balance rule operates in any given temporal or network circumstance is belied by these results.
What is also clear is the z-ties tend to be significantly suppressed in the face of <pn> 2-path relations. Again, it suggests neutrality is less an option many times in the face of these 2-path conditions. This general tendency, though, does not always coincide with the years during which this was true for the <z.pp> balance correlations in Figure 5.
Figure 7: PZ (and Twin ZP) Family of Balance Correlations
3.3.3 PZ (and Twin ZP) Family
Figure 7 is our first case where all members of the family are excluded from any of Heider’s original predictions. Thus, it is difficult to say whether any of these correlations should be theoretically balanced. What we do observe, though, is whether they appear to be empirically balanced. That is, are there any general patterns that show persistently positive (balanced) correlations, or perhaps persistently negative (unbalanced) correlations?
The answer is, clearly, “yes” to both questions. We find that alliance relations (positive ties) are less likely in the face of many <pz> (or <zp>) 2-path relations, suggesting that <p.pz> is an empirically unbalanced configuration, at least outside the LOA era. And, by contrast we find that a neutral stance against another country is more likely in the face of many <pz> (or <zp>) 2-path relations, suggesting that <z.pz> appears to be an empirically balanced configuration (again, outside the LOA years).
Figure 8: NN Family of Balance Correlations
3.3.4 NN Family
Figure 8 provides the balance correlations for the <nn> family. The first plot in this figure provides general support of the Heiderian balance proposition that “an enemy of an enemy is a friend” in that for over half of the years the <p.nn> correlation is positive and significant. In the post LOA period, since WWII, the <z.nn>, the <z.nn> correlation is almost always negative, and often significantly so, suggesting that neutral stances are perhaps discouraged with an enemy of an enemy.
Of interest in the <z.nn> and <n.nn> plots is the large bands areas of gray, which represent the range of correlations one is likely to observe if the null hypothesis were true. In particular, in the <n.nn> plot, the null region often ranges from a correlation of zero up to 1.0. While most of the observed correlations are either 0 or close to it, the wide range of possible null values suggests that we should not be confident in much of anything in this plot. This is largely because the density of n-ties is low to begin with, and thus the density of <nn> 2-path ties to be even less frequent. This scarcity permits the correlation to be quite unstable under the null hypothesis.
Figure 9: NZ (and Twin ZN) Family of Balance Correlations
3.3.5 NZ (and Twin ZN) Family
Figure 9 shows the results for the <nz> (or <zn> twin) family. Almost all of these balance correlations are either zero of close to zero, with the minor exception of the significantly positive <n.nz> correlations during the two world wars. Not only are the preponderance of these balance correlations near zero, they also are either just barely border the null region or clearly not significant at all. Thus, we have no evidence that a country who is an enemy of a neutral partner is likely to differentially engage in any particular relation (p, n or z) with another nation.
Figure 10: ZZ Family of Balance Correlations
3.3.6 ZZ Family
Our final set of results are provided in Figure 10. Z ties, and <zz> 2-path ties, are the most dense. Thus, as is displayed in the plots, the null regions are relatively thin. The <n.zz> balance correlations are still largely insignificant, either inside or bordering the null region.
It is worth deliberating on the meaning of these <zz> structures. These suggest that the Ego  Alter dyad is surrounded by non-ties or neutral ties to an inordinate number of other countries. It is as if Ego and Alter are in separate or separated worlds. Thus, a <p.zz> pattern would suggest that Ego and Alter tend to form a bridge between these relatively disconnected parties. A positive balance correlation for <p.zz> would indicate that an alliance is more likely if they are from these distant parts; a negative correlation would indicate that Ego and Alter are less likely to form an alliance if they are from these separated parts. What we see from the <p.zz> plot is that the balance correlations are generally if only slightly negative, suggesting alliances are discouraged across these chasms.
A <z.zz> triple tells a slightly different story. In this case, a positive correlation would indicate that Ego  Alter pairs across such gulfs are relatively more likely to stay neutral with each other. This would suggest that there are relatively disconnected clusters in the community, and that inside these clusters are where much of the action (both positive and negative) are happening. Indeed, the <z.zz> balance correlations, again outside the LOA period, are positive, if not exceptionally strong.
What is interesting in this family of plots, however, is that for the <p.zz> and <z.zz> balance correlations outside the LOA years, we find corelations that are significant (not contained in the null region) yet closer to zero than the null region itself. This pattern is particularly striking in the post LOA years, during and following WWII. The <p.zz> correlations are predominantly negative and yet they are significantly greater than we would expect by chance under the null.
4 Discussion and Extensions
This paper deals with signed graphs, where edges between nation states are undirected symmetric ties. One of Holland and Leinhardt’s major advances in this line of work was to recognize that networks are often, perhaps most often, composed of directed ties. The work presented here can be extended to directed ties, also, leading to new specific predictions and explorations of possible balance statements and commensurate balance correlations to assess them.
However, it should be noted that this increases substantially the number of possible balance statements. Instead of one relation for each pair, there is both an in-tie and an out-tie that would have to be specified; thus, for each pair we must specify two relations, not just one. For example, we could describe a case where Ego sent a positive tie to Alter, but Alter does not respond with any sentiment (i.e., send a z-tie back). Perhaps Ego sends a negative tie to X, and X reciprocates with a non-tie. Further, suppose Alter sends a negative tie to X and X responds with a similar negative tie back to Alter. This configuration would be described as <pz . nz nn>. Since there are six directed ties in any labeled triple, then the number of possible directed triad signatures is $7293^{6}=$ 729, not simply $27$ as described in this paper. An interesting extension of our approach to digraphs is that of these $729$ signatures, $702$ have an indistinguishable twin and $27$ do not (see the proof of the theorem 2.1 in the appendix). We expound on these structures, their balance statements, and their empirical results in a subsequent paper on balance correlations in directed graphs.
5 Conclusion
Our intent here is to propose a new approach to the theorizing and study of balance processes in network contexts. By formally incorporating neutral or zero ties into the analysis, we extend balance theory in a significant way that better matches the real world of actors in a systemic relational structure. And, in an important way, we return to Heider’s original thinking by focusing on Ego’s choice of how to relate to a specific Alter when confronted with the concommitant relations with one or many third parties. We present a novel nomenclature, the triad signature, that assists in identifying and theorizing about balanceimbalance conditions.
Finally, we propose a simple, intuitive measure, the balance correlation, that directly reflects the extent to which Heider’s predictions (and its extentions in this paper) are supported by the evidence in an observed network. This measure captures the difference in probabilities of Ego’s choice of a tie with Alter, conditioned on whether the specific balance-theoretic 2-path configuration exists or does not exist. We also propose a reasonable null test of this statistical measure and apply it to a well-known data set employed by political scientists in international affairs, the Correlates of War data.
As powerful as this approach is, we also note that the 27 different theoretical predictions that can be identified using the triad signature are not always uniquely identifiable empirically. We show that, indeed, 18 of these signature have a phenotypical twin signature, in that the two signatures will necessarily receive identical support in any given network and can therefore never be distinguished using a balance correlation. We also provide a proof that reveals the necessary and sufficient conditions for twins to exist and how to identify them.
We further demonstrate the power of this approach by exploring the balance correlations of all the possible predictions identified by the triad signatures over almost two centuries of conflict and alliances among extant nations in the world. We find systematic underlying patterns that deserve attention. First, we find strong evidence overall for balance theory on positive ties. In cases where the conditional 2-path relations contain either two positive ties or two negative ties, the probability of a direct positive tie is significant for most of the years, as would be expected according to balance theory.
Second, we find that the predictions around unbalanced states are not confirmed for the most part and occasionally even counter to Heider’s original contention. For example, the theoretically unbalanced <n.pp> correlation hovers around zero for most of the two centuries of data. And a clear example of a result that disconfirms an unbalanced prediction is given in Figure 6, where the theoretically unbalanced <p.pn> signature is consistently positive in the post WWII era.
Thirdly, in a related general finding, it appears that the z-ties (neutral ties) play a larger role in establishing balance, or at least avoiding unbalanced states, than negative hostile ties over substantial time periods. For example, in the case of where the Ego  Alter dyad is embedded in many <pp> 2-path relations (see Figure 5), there is do discernable evidence that Ego differentially avoids negative ties (<n.pp>), but there is substantial and significant suppression or avoidance of neutral ties (<z.pp>).
In sum, it is our contention that balance correlations provide a simple, intuitive way to assess the extent to which specific balance predictions might be true on the average across an entire network. Furthermore, by incorporating agentic zeroes into the exploration of balance conditions, we may uncover new properties of social order that have been prevalent, even dominant, but hidden from the researcher’s purvue or agenda. We hope this new approach opens up avenues of exploration that have been heretofore ignored.
Appendices
Appendix A Proof of theorem 2.1
The claim in Theorem 2.1 is based on the fact that if the indices of the two variables in the balance correlation are permuted in the same way, the correlation does not change due to associative and commutative properties of the addition of the cross-products in the correlation function. Transposing matrices is one unique permutation of the elements in a matrix.
Definition A.1.
Let $\mathfrak{M}=\{M_{1},...,M_{m}\}$ be the set of $m$ square matrices. All matrices $M_{q}$ are non-empty with zero diagonals and are binary such that each element $M_{q_{ij}}\in\{0,1\}$. Further, $m\geq 2$, and $\sum_{q=1}^{m}M_{q}=\mathfrak{L}$, is the matrix representation of the complete graph. Also, all matrices in $\mathfrak{M}$ have identical and identically ordered row and column indices, $i,j\in\{1,...,n\}$ for the $n$ number of nodes.
Definition A.2.
Let $\mathfrak{M}^{{}^{\prime}}=\{M^{{}^{\prime}}_{1},...,M^{{}^{\prime}}_{m}\}$ be the set of transposes of matrices in $\mathfrak{M}$. A nugget is defined to be the element-wise product of any matrix in $\mathfrak{M}$ with any matrix in $\mathfrak{M}^{{}^{\prime}}$. The nugget set $\mathfrak{U}=\mathfrak{M}\times\mathfrak{M}^{{}^{\prime}}=\{U_{1},...,U_{g}\}$ and contains all possible $g$ nuggets. A nugget thus has the same indices of the matrices in $\mathfrak{M}$ and is also binary. There are $g=m^{2}$ possible nuggets.
Lemma A.1.
In the nugget set $\mathfrak{U}$ consisting of $g$ nuggets, there will be exactly $m$ symmetric nuggets, and exactly $m^{2}-m$ asymmetric nuggets.
Proof of lemma A.1.
Note that $M_{q_{ji}}=M_{q_{ij}}^{{}^{\prime}}$. By contradiction, assume for any $ij$-th element in a nugget, the element-wise product of matrices, $M_{r_{ij}}\times M_{s_{ji}}=1$ and $M_{r_{ji}}\times M_{s_{ij}}=1$. Since, $M_{q_{ij}}\in\{0,1\}$, this implies $M_{r_{ij}}=M_{s_{ji}}=1$ and $M_{r_{ji}}=M_{s_{ij}}=1$, so $M_{r_{ij}}+M_{s_{ij}}=2$. However, if $r\neq s$ this contradicts that the sum over all matrices results in a matrix $\mathfrak{L}$ (see definition A.1). If and only if $r=s$ the assumption holds, as the summation visits each matrix $q$ in $\mathfrak{M}$ only once, which implies there are exactly $m$ symmetric nuggets; those of a matrix and it’s transpose. ∎
Definition A.3.
Let a balance correlation be the Pearson correlation between elements of a nugget $U_{k}$ and the elements of the inner product of two nuggets $U_{l}U_{m}$, ignoring the diagonals, i.e.,
$\rho(U_{k},U_{l}U_{m})=\frac{cov(U_{k},U_{l}U_{m})}{\sqrt{var(U_{k})var(U_{l}U_{m})}}$ (9)
Now, $\mathfrak{B}$, is the set containing all possibly unique balance correlations that can be derived from the $g$ nuggets in any set $\mathfrak{U}$. The specific order of the nuggets (and hence the matrices they are derived from) in a balance correlation, reveal a statement about the association between a 2-path and tie in bi-directional relations.
So, $\rho(U_{k},U_{l}U_{m})$ is a quantification of the statement about an association between ties in $U_{k}$ and 2-paths in $U_{l}U_{M}$. This quantification differs from $\rho(U_{k},U_{m}U_{l})$ or $\rho(U_{l},U_{k}U_{m})$, which quantify statements of association between ties in $U_{k}$ and $U_{m}U_{l}$, and, $U_{l}$ and $U_{k}U_{m}$, respectively.
Now, since the bi-variate Pearson correlation is defined in terms of variances and covariance, and those metrics are permutation invariant, i.e. for any permutation function $\pi_{p}$ we have
$\rho(U_{k},U_{l}U_{m})=\rho(\pi_{p}(U_{k}),\pi_{p}(U_{l})\pi_{p}(U_{m})).$ (10)
Given that there are $m^{2}$ nuggets, and, balance correlation considers 3 nuggets there are $m^{6}$ possible balance correlations. However these are not all unique, because transposition is a permutation and by definition A.2, nuggets include transposed matrices.
Nuggets, as Hadamard products that maintain associativity, imply that the order of matrices within a nugget is not important ($M_{r}\times M_{s}^{{}^{\prime}}=M_{s}^{{}^{\prime}}\times M_{r})$. However, it is relevant which matrix in a nugget is transposed in revealing the underlying theoretical statement. Furthermore, the second variable, $U_{l}U_{m}$, is a matrix inner product. It is well-known that taking the transpose of an inner product, transposes the matrices and reverses their order. There are three situations in which transposing either or both variables in $\rho$ changes the formulation of the underlying statement or in other words demonstrate equivalence between two underlying statements under all possible instantiates. Only, when both $U_{k}$ and $U_{l}U_{m}$ are symmetric by construction $\rho$ quantifies a sole statement. In the other cases we speak of twin statements.
Proof of theorem 2.1.
Assume nuggets $\{U_{k},U_{l},U_{m}\}\in\mathfrak{U}$ and define $\{U_{k}=U_{q}^{{}^{\prime}},U_{l}=U_{r}^{{}^{\prime}},U_{m}=U_{s}^{{}^{\prime}}\}$ which implies $\{U_{q},U_{r},U_{s}\}\in\mathfrak{U}$. Since, transposing is a permutation, by equation 10 we have
$\rho(U_{k},U_{l}U_{m})=\rho(U_{k}^{{}^{\prime}},U_{m}^{{}^{\prime}}U_{l}^{{}^{\prime}}),$ (11)
which by substitution gives $\rho(U_{k},U_{l}U_{m})=\rho(U_{q},U_{s}U_{r})$. If and only if, $U_{k}=U_{k}^{{}^{\prime}}\neq U_{q}$, and, $U_{l}U_{m}=U_{m}^{{}^{\prime}}U_{l}^{{}^{\prime}}\neq U_{s}U_{r}$ no other quantification exists that will always produce the same quantity. By the fact that there exists one and only one possible transpose of each nugget, in all other cases there exists exactly one other configuration of distinguishable nuggets that will lead to the same quantification, where $U_{k}=U_{k}^{{}^{\prime}}=U_{q}$, and, $U_{l}U_{m}=U_{m}^{{}^{\prime}}U_{l}^{{}^{\prime}}=U_{s}U_{r}$. ∎
It follows that quantifying a sole statement requires a symmetric nugget $U_{k}$, which implies by lemma A.1 the matrices in $U_{k}$ are $M_{t}$ and $M_{t}^{{}^{\prime}}$, as well as that $U_{l}=U_{m}^{{}^{\prime}}$, which implies $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{w}M_{v}^{{}^{\prime}}$. In those cases, one and only one configuration of nuggets (and matrices) exists, since transposing equals the trivial identity permutation. If $U_{k}=U_{q}^{{}^{\prime}}$ for any $U_{k}=M_{t}\times M_{s}^{{}^{\prime}}$ and $U_{q}=M_{s}\times M_{t}^{{}^{\prime}}$ will contain the same set of values, but in a different order. Similarly, in any $U_{l}U_{m}$ that has $U_{s}U_{r}$ as a distinguishable transpose, all values $U_{l}U_{m}$ uniquely exist as transpose in $U_{s}U_{r}$, and vice verse. So for any $U_{l}=M_{v}M_{w}^{{}^{\prime}}$ and $U_{m}=M_{x}M_{y}^{{}^{\prime}}$ there exists $U_{r}=M_{w}M_{v}^{{}^{\prime}}$ and $U_{s}=M_{y}M_{x}^{{}^{\prime}}$, such that $\sum_{k}(M_{v_{ik}}M_{w_{ki}})(M_{x_{kj}}M_{y_{jk}})=\sum_{k}(M_{y_{jk}}M_{x_{kj}})(M_{v_{ik}}M_{w_{ki}})\,\forall\,\{i,j,k\}$. Hence, only when either or both variables in $\rho$ have a distinguishable transpose a distinguishable twin statement exists that produces the exact same value for $\rho$, i.e. showing duality for this subset of balance correlations.
Cardinality
Definition A.3 determines any balance correlation $\rho(U_{k},U_{l}U_{m})$ is a bi-variate Pearson correlation. This leads us to consider two conditions, or distinguishing assumptions, as special cases. Although we recognise more complex situations could apply, these two limiting conditions will provide an upper and lower cardinality bound for $\mathfrak{B}$.
Assumption A.1.
All matrices in $\mathfrak{M}$ are symmetric.
This implies that for all nuggets in =$\mathfrak{Z}=\{M_{r}\times M_{s}^{{}^{\prime}}\,|\,r\neq s,M_{r}\in\mathfrak{M},M_{s}^{{}^{\prime}}\in\mathfrak{M}^{{}^{\prime}}\}$ the variances $var(Z_{h})=0$. Since the Pearson correlation is defined if and only if $var(U_{k})>0$ and $var(U_{l}U_{m})>0$, the balance correlations that include $Z_{h}$ are not defined, and cannot be considered. Therefore, under assumption A.1 the nugget set cardinality will be $|\mathfrak{U}|=m$, and will only contain symmetric nuggets.
The symmetry of the nuggets implies there are $m$ possible nuggets for the first variable in $\rho$. Because, there is no restriction on the use of nuggets in any of the $2$ positions in the second variable, the number of possible combinations to constitute it, in general, is given by
$\eta=\Biggl{(}\binom{m}{2}\Biggr{)}=\frac{(m-1+2)!}{(m-1)!2!},$ (12)
where $m$ is the number of different nuggets in $\mathfrak{Z}$. It follows that $|\mathfrak{B}|=\eta\cdot m$ is the cardinality of the largest possible set of unique balance correlations.
In the specific case discussed in this paper, where we are limiting ourselves to a set of three symmetric relations (P, N, and Z), $\eta=6$ and thus the largest possible set of unique correlations $=3\cdot 6=18$. Therefore, of the $m^{3}=27$ possible balance correlations, only 18 of them will be unique, and the remaining nine will be twins, paired with nine elements in this set of 18 unique balance correlations. With nine pairs of twins, that means the remaining nine balance correlations will have no twins.
Assumption A.2.
All matrices in $\mathfrak{M}$ are asymmetric, nuggets have $var(U_{k})>0\,\forall\,k\in\{1,...,g\}$, and the inner product of nuggets have $var(U_{l}U_{m})>0\,\forall\,l,m\in\{1,...,g\}$.
The proof of theorem 2.1 provides the conditions under which a specific configuration of nuggets doesn’t have a twin. It requires a symmetric nugget in the first variable in $\rho$, and a symmetric inner product of nuggets in the second variable. The latter condition can only hold when no more than two different matrices constitute the two nuggets. The order of the two matrices do define the specific nuggets, and hence is important. However, these two nuggets also need to be transposes of each other. Given $m$ asymmetric matrices there are exactly $m$ symmetric nuggets, according to the proof of lemma A.1. Furthermore, there are exactly $g=m^{2}$ nuggets, which includes their transposes. So, the total number of possible inner products equals $m^{4}$, of which $m^{2}$ will be inner products with the transpose. Hence, there are $\alpha=m\cdot m^{2}=m^{3}$ number of configurations of matrices in the variables of $\rho$ that quantify a sole statement. The other possible balance correlations all have exactly one other equal quantification of a twin statement. Therefore,
$|\mathfrak{B}|=\frac{1}{2}(m^{6}-m^{3})+m^{3}=\frac{1}{2}m^{3}(m^{3}+1)$ (13)
where $m^{6}$ is the number of combinations of all possible nuggets (2 matrices in a nugget) in the three nuggets used in $\rho(U_{k},U_{l}U_{m})$. For any set of mixed symmetric and asymmetric matrices the cardinality of $|\mathfrak{B}|$ will be between those in equations 12 and 13.
The logical entanglement of statements is both interesting and limiting this approach. In fact it can be solved by extending the approach to consider an extra temporal dimension to the data. Note however, that the balanced states all rely on the assumption of reciprocal relations (symmetric nuggets). As such, at any moment in time these states, according to the theory, may be expected to occur more than unbalanced states. This motivates the use of balance correlation as a cross-sectional concept, and to rely on a subset of $|\mathfrak{B}|$, which reflects the statement. Yet, from an exploratory perspective it remains of interest to inspect the values of other measures as this can reveal information about the process of balance formation.